OpenStudy (anonymous):

Teri got a 44 and 72 on her 1st 2 algebra tests. In order to pass the class with a 60%, what is the lowest grade she can get on the remaining test????

OpenStudy (mathteacher1729):

Assuming that the entire class grade is just the average of the three tests (a poorly designed grading scheme if you ask me!) then you have to solve the following: \[\frac{44+72+x}{300}=\frac{60}{100}\] This is not so bad because you can add 44 and 72 because they are in the numerator of the left fraction... \[\frac{116+x}{300}=\frac{60}{100}\] Now reduce that 60/100 on the right hand side: Divide by 10 \[\frac{116+x}{300}=\frac{6}{10}\] Divide by 2 \[\frac{116+x}{300}=\frac{3}{5}\] Ok, so from here we do some cross multiplication to solve for x. :) Hope this helps!